# appendixtocaptAppendix

### v3

@@ D.2 Properties of the Gradient Field
$p(x|y,t)=\frac{1}{4\pi\alpha t}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha t}\right),\quad\alpha=1/4$ (53)
Since $p(x|x,t)=p(y|y,t)=\frac{1}{4\pi\alpha t}$, the normalization gives:
$q(x|y,t)=\exp\left(-\frac{\|x-y\|^{2}}{4\alpha t}\right)$ (54)
@@
$\nabla_{x}q(x|y,t)=-\frac{x-y}{2\alpha t}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha t}\right)\propto-(x-y)$ (55)
Thus, $\frac{dx(\tau)}{d\tau}=\nabla_{x}q(x|y,t)$ follows a straight line from $x$ to $y$, matching the Euclidean shortest path [6].
For small $t$, $q(x|y,t)$ aligns with the geodesic distance:
$p(x|y,t)\sim\frac{1}{(4\pi\alpha t)^{d/2}}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha t}\right)$ (56)
where $d_{g}(x,y)$ is the geodesic distance [60]. Since $q(x|y,t)\propto p(x|y,t)$ in an open field, $\nabla_{x}q(x|y,t)\propto-\nabla_{x}d_{g}^{2}(x,y)$, ensuring the planned path follows the shortest route, even around obstacles.
For large $t$, $q(x|y,t)$ emphasizes global topological connectivity over local geometric details, often producing paths superior to the shortest. The transition matrix $P_{t}=P_{1}^{t}$ has spectral decomposition:
$P_{t}=Q\Lambda^{t}Q^{T}$ (57)
where $Q$ is orthogonal, and $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$ contains eigenvalues $0\leq\lambda_{i}\leq 1$ [10]. For large $t$, dominant eigenvalues ($\lambda_{i}\approx 1$) amplify global connectivity:
$p(x|y,t)=[P_{t}]_{xy}\approx\sum_{i:\lambda_{i}\approx 1}\lambda_{i}^{t}Q_{i}(x)Q_{i}(y)$ (58)
The eigenvectors $Q_{i}$ corresponding to $\lambda_{i}\approx 1$ encode topological features, such as major pathways and passage connectivity, invariant to deformations like stretching [10]. Thus, $\nabla_{x}q(x|y,t)$ prioritizes routes with high connectivity (e.g., wider corridors or multiple paths), potentially deviating from the shortest path to favor robust, flexible trajectories [44]. For example, in a maze, the planned path may choose a longer but more reliable route through a well-connected passage, avoiding narrow or risky shortcuts.
@@
The planned path’s alignment with the shortest path for small $t$ and its topological robustness for large $t$ mirror the hippocampus’s ability to balance precision and global structure, ensuring efficient navigation across diverse environments.
D.4 Boundary Layer Effects
Near obstacles, the gradient field $\nabla_{x}q(x|y,t)$ aligns parallel to boundaries, preventing collisions. This behavior arises from the reflecting boundary condition of the heat equation governing $p(x|y,t)$, with ${\partial p(x|y,t)}/{\partial n}=0$ on $\partial\Omega_{\text{obstacles}}$. The condition ensures the normal component of $\nabla_{x}p(x|y,t)$ vanishes at boundaries, creating a flow parallel to obstacles, with strength modulated by the goal’s position and time scale $t$.
D.5 Diffraction-Like Patterns
Diffraction, in the context of our navigation framework, refers to the bending of the gradient field $\nabla_{x}q(x|y,t)$ around obstacles, such as corners or narrow passages, analogous to how waves bend around edges in optics or acoustics. This phenomenon arises because the transition probability $p(x|y,t)$, governed by the heat equation, diffuses probability mass through available paths, concentrating gradients toward openings like passages or around corners. These diffraction-like patterns guide trajectories through high-gradient regions, ensuring efficient navigation in complex environments.
@@ D.5 Diffraction-Like Patterns
$p(x|y,t+1)=\sum_{z}p(x|z,t)\,p(z|y,1)$ (59)
This convolution reflects the diffusion of probability from $y$ to $x$ through intermediate points $z$, allowing $p(x|y,t)$ to “bend” around obstacles by accumulating contributions from multiple paths [24]. Near a corner or passage, the heat kernel explores paths that wrap around the obstacle, creating a gradient field that converges toward navigable routes.
For a narrow passage of width $w$, the transition probability approximates the geodesic distance $d_{g}(x,y)$ through the passage. For small to moderate $t$:
$q(x|y,t)\approx\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha t}\right)$ (60)
@@
$\nabla_{x}q(x|y,t)\approx-\frac{1}{2\alpha t}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha t}\right)\nabla_{x}d_{g}^{2}(x,y)$ (61)
Since $\nabla_{x}d_{g}^{2}(x,y)$ points along the geodesic path through the passage, $\nabla_{x}q(x|y,t)$ converges toward the opening, creating a funneling effect. The width of this convergence region scales with $\sqrt{t}$, as the heat kernel’s diffusion spreads over a distance proportional to $\sqrt{4\alpha t}$. For small $t$, the gradient tightly focuses on the passage, ensuring precise navigation. For large $t$, the broader diffusion enables early detection of passages from greater distances, smoothing trajectories by integrating multiple nearby paths.
At sharp corners, diffraction-like patterns emerge similarly. The heat kernel accumulates probability along paths bending around the corner, creating a curved gradient field that guides the agent past the obstacle. Unlike passages, corners lack a single geodesic path, so the gradient field spreads, resembling optical diffraction patterns where intensity peaks near edges. This spreading is more pronounced for large $t$, as the heat kernel explores longer, circuitous routes.
These patterns ensure robust navigation in complex environments. The funneling effect through passages and curved trajectories around corners mirror hippocampal navigation, where place cells encode paths that navigate mazes or avoid obstacles [38, 41]. The $t$-dependent scaling reflects the hippocampus’s ability to adjust focus, balancing precision for small $t$ with global awareness for large $t$, enhancing adaptability in varied spatial contexts [60].
D.6 Topological Invariance
Topological invariance ensures that navigation behavior, driven by the gradient field $\nabla_{x}q(x|y,t)$, remains consistent under continuous deformations (e.g., stretching, bending) that preserve the environment’s connectivity, such as the number of passages or loops. This robustness mirrors the hippocampus’s cognitive maps, which encode abstract spatial relationships invariant to physical distortions [59].
The transition probability $p(x|y,t)$ arises from a random walk on a discrete grid, with one-step transition matrix $P_{1}$. For time $t$, the transition matrix is:
$P_{t}=P_{1}^{t}$ (62)
@@ D.6 Topological Invariance
For large $t$, dominant eigenvalues ($\lambda_{i}\approx 1$) amplify global topological features:
$P_{t}\approx\sum_{i:\lambda_{i}\approx 1}\lambda_{i}^{t}Q_{i}Q_{i}^{T}$ (64)
@@
The gradient field is $\nabla_{x}q(x|y,t)$, where $p(x|y,t)=[P_{t}]_{xy}$. Since $P_{t}$’s eigenvectors are topologically invariant, $\nabla_{x}q(x|y,t)$ preserves navigation paths (e.g., through passages) across equivalent environments. For large $t$, the dominance of $\lambda_{i}\approx 1$ ensures that $\nabla_{x}q(x|y,t)$ prioritizes global connectivity, guiding trajectories along major routes regardless of local geometry.
This invariance aligns with hippocampal navigation, where place cells encode connectivity (e.g., room layouts) over precise metrics, enabling robust path planning under environmental changes [41]. The focus on topology for large $t$, driven by eigenvectors tied to $\lambda_{i}\approx 1$, reflects cognitive maps that emphasize structural relationships, as observed in rodent navigation [38, 59].
D.7 Hippocampal Preplay and Shortcut Detection
@@
The one-step transition matrix $P_{1}$ defines probabilities of moving between adjacent grid points. Multi-step transition matrices $P_{t}=P_{1}^{t}$ for scales $t=2^{k}$ ($k=1,2,\ldots,K$) are computed via matrix squaring:
@@
The path planning algorithm adaptively selects $t^{*}=\operatorname*{arg\,max}_{t}\|\nabla_{x}q(x|y,t)\|$, guiding trajectories via $\nabla_{x}q(x|y,t^{*})$. For regions $A$ and $B$ connected by an unexplored passage, local exploration defines $P_{1}$. Matrix squaring yields $P_{t}$, and adaptive $t$-selection ensures $q(a|b,t^{*})>0$, discovering the shortcut without prior traversal.
@@
This process—matrix squaring to compute $P_{t}$ and adaptive $t$-selection—embodies hippocampal preplay, predicting connectivity from local data, akin to place cell sequence pre-activation. The framework’s efficiency and reliance on $P_{1}$ enhance its biological plausibility, supporting robust navigation.
Appendix E Theta Phase Modeling
@@
We propose a local definition of theta phase that embeds spatial adjacency within the phase angle, capturing the observed precession range of 270° to 90° while enabling navigation across a tessellated environment. This formulation leverages an external  8 Hz theta rhythm for temporal pacing, where $\tau$ denotes real time during navigation, distinct from the random walk time scale $t$, and aligns with distributed hippocampal computation without requiring individual cells to explicitly encode spatial centers.
E.1 Local Phase from Embedding Similarity
@@
Define the adjacency of the current position $x(\tau)$ to the field center of the $i$-th place cell at scale $t$:
$a_{i}(\tau)=\langle h(x(\tau),t),h(\mu_{i},t)\rangle$ (66)
@@
where $h(x(\tau),t)\in\mathbb{R}^{n}$ is the population embedding (firing rates of $n$ place cells) at position $x(\tau)$, and $h(\mu_{i},t)$ represents the embedding at the center $\mu_{i}$, where
$h_{i}(\mu_{i},t)=\max_{x\in\mathcal{X}}h_{i}(x,t)$ (67)
Collectively $(\mu_{i},\forall i)$ form a tiling or discretization of the environment.
@@
We interpret $h(\mu_{i},t)=[w_{1i},w_{2i},\ldots,w_{ni}]$ as the recurrent connection weights from the population to neuron $i$, such that:
$a_{i}(\tau)=\sum_{j=1}^{n}h_{j}(x(\tau),t)w_{ji}$ (68)
@@
representing the net recurrent input to neuron $i$. This corresponds to the normalized transition probability $q(x(\tau)|\mu_{i},t)$, and in open fields (Appendix C), it follows a Gaussian profile:
$a_{i}(\tau)=\exp\left(-\frac{\|x(\tau)-\mu_{i}\|^{2}}{t}\right)$ (69)
The theta phase is defined based on $\arccos(a_{i}(\tau))$, which is the angle between $h(x(\tau),t)$ and $h(\mu_{i},t)$:
$\phi_{i}(x(\tau),t)=\pi+\text{sgn}\left(\frac{da_{i}(\tau)}{d\tau}\right)\cdot\left(\frac{\pi}{2}-\arccos(a_{i}(\tau))\right)$ (70)
@@
where: $\frac{da_{i}(\tau)}{d\tau}=\sum_{j=1}^{n}\frac{dh_{j}(x(\tau),t)}{d\tau}w_{ji}$ is the temporal rate of change of this recurrent input, reflecting movement toward ($>0$) or away ($<0$) from $\mu_{i}$ as encoded by the weights. The angle $\arccos(a_{i}(\tau))\in[0,\pi/2]$ embeds $a_{i}(\tau)$ into the phase, possibly by sinusoidal inhibitory mechanism [30, 26]. This angle-phase duality is elegant and compelling.
@@
    At field entry ($a_{i}\to 0$, $\frac{da_{i}}{d\tau}>0$): $\phi_{i}=\frac{3\pi}{2}$ (270°), late phase.
@@
    At field center ($a_{i}=1$, $\frac{da_{i}}{d\tau}=0$): $\phi_{i}=\pi$ (180°), mid-phase.
@@
    At field exit ($a_{i}\to 0$, $\frac{da_{i}}{d\tau}<0$): $\phi_{i}=\frac{\pi}{2}$ (90°), early phase.
This does not require neuron $i$ to encode $\mu_{i}$ explicitly; rather, $h(\mu_{i},t)$ emerges as a distributed representation within recurrent connectivity, learned through experience and computed collectively, with $\phi_{i}$ storing $a_{i}(\tau)$ as an angular signal.
E.2 Implications for Navigation
This phase also enables navigation by tessellating the environment with $\{\mu_{i}\}$. Each $\mu_{i}$ anchors a tile, and $\phi_{i}$ signals adjacency: 180° indicates proximity to $\mu_{i}$, while 270° and 90° mark entry and exit. For a goal at $\mu_{g}$, navigating toward $\phi_{g}=\pi$ (max $a_{g}$) guides movement, with $\{\phi_{i}\}$ identifying intermediate $\mu_{i}$ to approach $\mu_{g}$. This discretizes continuous navigation into a phase-driven sequence over $\{\mu_{i}\}$, complementing our gradient-based approach with a recurrently-driven strategy. During rest, iterating $\phi_{i}$ at  8 Hz drives replay, with $\frac{da_{i}}{d\tau}$ reflecting virtual transitions, unifying coding and navigation.

### v4

@@ G.2 Properties of the Gradient Field
$p(x|y,\tau)=\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha\tau}\right),\quad\alpha=1/4$ (64)
Since $p(x|x,\tau)=p(y|y,\tau)=\frac{1}{4\pi\alpha\tau}$, the normalization gives:
$q(x|y,\tau)=\exp\left(-\frac{\|x-y\|^{2}}{4\alpha\tau}\right)$ (65)
@@
$\nabla_{x}q(x|y,\tau)=-\frac{x-y}{2\alpha\tau}\exp\left(-\frac{\|x-y\|^{2}}{4\alpha\tau}\right)\propto-(x-y)$ (66)
Thus, $\frac{dx(t)}{dt}=\nabla_{x}q(x|y,\tau)$ follows a straight line from $x$ to $y$, matching the Euclidean shortest path [64].
For small $\tau$, $q(x|y,\tau)$ aligns with the geodesic distance:
$p(x|y,\tau)\sim\frac{1}{(4\pi\alpha\tau)^{d/2}}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha\tau}\right)$ (67)
where $d_{g}(x,y)$ is the geodesic distance [16]. Since $q(x|y,\tau)\propto p(x|y,\tau)$ in an open field, $\nabla_{x}q(x|y,\tau)\propto-\nabla_{x}d_{g}^{2}(x,y)$, ensuring the planned path follows the shortest route, even around obstacles.
For large $\tau$, $q(x|y,\tau)$ emphasizes global topological connectivity over local geometric details, often producing paths superior to the shortest. The transition matrix $P_{\tau}=P_{1}^{\tau}$ has spectral decomposition:
$P_{\tau}=Q\Lambda^{\tau}Q^{T}$ (68)
where $Q$ is orthogonal, and $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$ contains eigenvalues $0\leq\lambda_{i}\leq 1$ [57]. For large $\tau$, dominant eigenvalues ($\lambda_{i}\approx 1$) amplify global connectivity:
$p(x|y,\tau)=[P_{\tau}]_{xy}\approx\sum_{i:\lambda_{i}\approx 1}\lambda_{i}^{\tau}Q_{i}(x)Q_{i}(y)$ (69)
The eigenvectors $Q_{i}$ corresponding to $\lambda_{i}\approx 1$ encode topological features, such as major pathways and passage connectivity, invariant to deformations like stretching [57]. Thus, $\nabla_{x}q(x|y,\tau)$ prioritizes routes with high connectivity (e.g., wider corridors or multiple paths), potentially deviating from the shortest path to favor robust, flexible trajectories [10]. For example, in a maze, the planned path may choose a longer but more reliable route through a well-connected passage, avoiding narrow or risky shortcuts.
@@
The planned path’s alignment with the shortest path for small $\tau$ and its topological robustness for large $\tau$ mirror the hippocampus’s ability to balance precision and global structure, ensuring efficient navigation across diverse environments.
G.4 Boundary Layer Effects
Near obstacles, the gradient field $\nabla_{x}q(x|y,\tau)$ aligns parallel to boundaries, preventing collisions. This behavior arises from the reflecting boundary condition of the heat equation governing $p(x|y,\tau)$, with ${\partial p(x|y,\tau)}/{\partial n}=0$ on $\partial\Omega_{\text{obstacles}}$. The condition ensures the normal component of $\nabla_{x}p(x|y,\tau)$ vanishes at boundaries, creating a flow parallel to obstacles, with strength modulated by the goal’s position and time scale $\tau$.
G.5 Diffraction-Like Patterns
Diffraction, in the context of our navigation framework, refers to the bending of the gradient field $\nabla_{x}q(x|y,\tau)$ around obstacles, such as corners or narrow passages, analogous to how waves bend around edges in optics or acoustics. This phenomenon arises because the transition probability $p(x|y,\tau)$, governed by the heat equation, diffuses probability mass through available paths, concentrating gradients toward openings like passages or around corners. These diffraction-like patterns guide trajectories through high-gradient regions, ensuring efficient navigation in complex environments.
@@ G.5 Diffraction-Like Patterns
$p(x|y,\tau+1)=\sum_{z}p(x|z,\tau)\,p(z|y,1)$ (70)
This convolution reflects the diffusion of probability from $y$ to $x$ through intermediate points $z$, allowing $p(x|y,\tau)$ to “bend” around obstacles by accumulating contributions from multiple paths [15]. Near a corner or passage, the heat kernel explores paths that wrap around the obstacle, creating a gradient field that converges toward navigable routes.
For a narrow passage of width $w$, the transition probability approximates the geodesic distance $d_{g}(x,y)$ through the passage. For small to moderate $\tau$:
$q(x|y,\tau)\approx\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha\tau}\right)$ (71)
@@
$\nabla_{x}q(x|y,\tau)\approx-\frac{1}{2\alpha\tau}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha\tau}\right)\nabla_{x}d_{g}^{2}(x,y)$ (72)
Since $\nabla_{x}d_{g}^{2}(x,y)$ points along the geodesic path through the passage, $\nabla_{x}q(x|y,\tau)$ converges toward the opening, creating a funneling effect. The width of this convergence region scales with $\sqrt{\tau}$, as the heat kernel’s diffusion spreads over a distance proportional to $\sqrt{4\alpha\tau}$. For small $\tau$, the gradient tightly focuses on the passage, ensuring precise navigation. For large $\tau$, the broader diffusion enables early detection of passages from greater distances, smoothing trajectories by integrating multiple nearby paths.
At sharp corners, diffraction-like patterns emerge similarly. The heat kernel accumulates probability along paths bending around the corner, creating a curved gradient field that guides the agent past the obstacle. Unlike passages, corners lack a single geodesic path, so the gradient field spreads, resembling optical diffraction patterns where intensity peaks near edges. This spreading is more pronounced for large $\tau$, as the heat kernel explores longer, circuitous routes.
These patterns ensure robust navigation in complex environments. The funneling effect through passages and curved trajectories around corners mirror hippocampal navigation, where place cells encode paths that navigate mazes or avoid obstacles [11, 9]. The $\tau$-dependent scaling reflects the hippocampus’s ability to adjust focus, balancing precision for small $\tau$ with global awareness for large $\tau$, enhancing adaptability in varied spatial contexts [16].
G.6 Topological Invariance
Topological invariance ensures that navigation behavior, driven by the gradient field $\nabla_{x}q(x|y,\tau)$, remains consistent under continuous deformations (e.g., stretching, bending) that preserve the environment’s connectivity, such as the number of passages or loops. This robustness mirrors the hippocampus’s cognitive maps, which encode abstract spatial relationships invariant to physical distortions [3].
The transition probability $p(x|y,\tau)$ arises from a random walk on a discrete grid, with one-step transition matrix $P_{1}$. For time $\tau$, the transition matrix is:
$P_{\tau}=P_{1}^{\tau}$ (73)
@@ G.6 Topological Invariance
For large $\tau$, dominant eigenvalues ($\lambda_{i}\approx 1$) amplify global topological features:
$P_{\tau}\approx\sum_{i:\lambda_{i}\approx 1}\lambda_{i}^{\tau}Q_{i}Q_{i}^{T}$ (75)
@@
The gradient field is $\nabla_{x}q(x|y,\tau)$, where $p(x|y,\tau)=[P_{\tau}]_{xy}$. Since $P_{\tau}$’s eigenvectors are topologically invariant, $\nabla_{x}q(x|y,\tau)$ preserves navigation paths (e.g., through passages) across equivalent environments. For large $\tau$, the dominance of $\lambda_{i}\approx 1$ ensures that $\nabla_{x}q(x|y,\tau)$ prioritizes global connectivity, guiding trajectories along major routes regardless of local geometry.
This invariance aligns with hippocampal navigation, where place cells encode connectivity (e.g., room layouts) over precise metrics, enabling robust path planning under environmental changes [9]. The focus on topology for large $\tau$, driven by eigenvectors tied to $\lambda_{i}\approx 1$, reflects cognitive maps that emphasize structural relationships, as observed in rodent navigation [11, 3].
G.7 Hippocampal Preplay and Shortcut Detection
@@
The one-step transition matrix $P_{1}$ defines probabilities of moving between adjacent grid points. Multi-step transition matrices $P_{\tau}=P_{1}^{\tau}$ for scales $\tau=2^{k}$ ($k=1,2,\ldots,K$) are computed via matrix squaring:
@@
The path planning algorithm adaptively selects $\tau^{*}=\operatorname*{arg\,max}_{t}\|\nabla_{x}q(x|y,\tau)\|$, guiding trajectories via $\nabla_{x}q(x|y,\tau^{*})$. For regions $A$ and $B$ connected by an unexplored passage, local exploration defines $P_{1}$. Matrix squaring yields $P_{\tau}$, and adaptive $\tau$-selection ensures $q(a|b,\tau^{*})>0$, discovering the shortcut without prior traversal.
@@
This process—matrix squaring to compute $P_{\tau}$ and adaptive $\tau$-selection—embodies hippocampal preplay, predicting connectivity from local data, akin to place cell sequence pre-activation. The framework’s efficiency and reliance on $P_{1}$ enhance its biological plausibility, supporting robust navigation.
Appendix H Theta Phase Modeling
@@
We propose a local definition of theta phase that embeds spatial adjacency within the phase angle, capturing the observed precession range of 270° to 90° while enabling navigation across a tessellated environment. This formulation leverages an external  8 Hz theta rhythm for temporal pacing, where $t$ denotes real time during navigation, distinct from the random walk time scale $\tau$, and aligns with distributed hippocampal computation without requiring individual cells to explicitly encode spatial centers.
H.1 From Inner Product Angle to Theta Phase
The emergent sparsity from non-negativity and orthogonality constraints endows each cell in $h(x,\tau)$ a localized place field, so that each cell $i$ has a place field centered at $\mu_{i}$. The place fields of all the cells tile the whole environment. At the current position $x(t)$ at real time $t$, the adjacency between $x(t)$ and each $\mu_{i}$ is measured by the angle between $h(x(t),\tau)$ and $h(\mu_{i},\tau)$, and this angle can then be embedded as the theta phase.
@@ H.1 From Inner Product Angle to Theta Phase
Specifically, define the adjacency of the current position $x(t)$ to the field center of the $i$-th place cell at scale $\tau$:
$a_{i}(t)=\langle h(x(t),\tau),h(\mu_{i},\tau)\rangle$ (77)
@@
where $h(x(t),\tau)\in\mathbb{R}^{n}$ is the population embedding (firing rates of $n$ place cells) at position $x(t)$, and $h(\mu_{i},\tau)$ represents the embedding at the center $\mu_{i}$, where
$h_{i}(\mu_{i},\tau)=\max_{x\in\mathcal{X}}h_{i}(x,\tau)$ (78)
@@
We interpret $h(\mu_{i},\tau)=[w_{1i},w_{2i},\ldots,w_{ni}]$ as the recurrent connection weights from the population to neuron $i$, such that:
$a_{i}(t)=\sum_{j=1}^{n}h_{j}(x(t),\tau)w_{ji}$ (79)
@@
representing the net recurrent input to neuron $i$. This corresponds to the normalized transition probability $q(x(t)|\mu_{i},\tau)$, and in open fields (Appendix E), it follows a Gaussian profile:
$a_{i}(t)=\exp\left(-\frac{\|x(t)-\mu_{i}\|^{2}}{\tau}\right)$ (80)
The theta phase is defined as:
$\phi_{i}(x(t),\tau)=\pi+\text{sgn}\left(\frac{da_{i}(t)}{dt}\right)\cdot\left(\frac{\pi}{2}-\arccos(a_{i}(t))\right)$ (81)
@@
where: $\frac{da_{i}(t)}{dt}=\sum_{j=1}^{n}\frac{dh_{j}(x(t),\tau)}{dt}w_{ji}$ is the temporal rate of change of this recurrent input, reflecting movement toward (>0>0) or away (<0<0) from $\mu_{i}$ as encoded by the weights. $\arccos(a_{i}(t))\in[0,\pi/2]$ is the angle between $h(x(t),\tau)$ and $h(\mu_{i},\tau)$, and this angle is embedded into the phase, possibly by sinusoidal inhibitory mechanism [65, 66].
@@
    At field entry ($a_{i}\to 0$, $\frac{da_{i}}{dt}>0$): $\phi_{i}=\frac{3\pi}{2}$ (270°), late phase.
@@
    At field center ($a_{i}=1$, $\frac{da_{i}}{dt}=0$): $\phi_{i}=\pi$ (180°), mid-phase.
@@
    At field exit ($a_{i}\to 0$, $\frac{da_{i}}{dt}<0$): $\phi_{i}=\frac{\pi}{2}$ (90°), early phase.
This does not require neuron $i$ to encode $\mu_{i}$ explicitly; rather, $h(\mu_{i},\tau)$ emerges as a distributed representation within recurrent connectivity, learned through experience and computed collectively, with $\phi_{i}$ storing $a_{i}(t)$ as an angular signal.
H.2 Implications for Navigation
This phase also enables navigation by tessellating the environment with $\{\mu_{i}\}$. Each $\mu_{i}$ anchors a tile, and $\phi_{i}$ signals adjacency: 180° indicates proximity to $\mu_{i}$, while 270° and 90° mark entry and exit. For a goal at $\mu_{g}$, navigating toward $\phi_{g}=\pi$ (max $a_{g}$) guides movement, with $\{\phi_{i}\}$ identifying intermediate $\mu_{i}$ to approach $\mu_{g}$. This discretizes continuous navigation into a phase-driven sequence over $\{\mu_{i}\}$, complementing our gradient-based approach with a recurrently-driven strategy. During rest, iterating $\phi_{i}$ at  8 Hz drives replay, with $\frac{da_{i}}{dt}$ reflecting virtual transitions, unifying coding and navigation.